Sphere A, with a charge of +6.0 ๐œ‡๐ถ, is located near another charged sphere, B. Sphere B has a charge of โˆ’3.0 ๐œ‡๐ถand is located 4.0 cm to the right of A.





What is the force of sphere B on sphere A?





A third sphere, C with a +1.5 ๐œ‡๐ถcharge, is added to the configuration. If it is located 3.0 cm directly beneath A, what is the new net force on sphere A?

1
Expert's answer
2022-03-08T08:56:24-0500

The Coulomb's law says


"F=k\frac{q_1q_2}{r^2}""F_1=9*10^9*\frac{6*10^{-6}*3*10^{-6}}{(0.04)^2}=101\:\rm N"